OpenStudy (matheducatormcg):

p is an odd prime. If p | x^p+y^p, then p^2 | x^p+y^p. Any helpful hints to prove this? I know it involves a contradiction with Fermat's Little Theorem.

OpenStudy (matheducatormcg):

Jumping into a meeting, but am eager to see if you have any insight #Zarkon.

OpenStudy (zarkon):

I have an idea...it uses Fermat's little theorem, but it is a direct proof

OpenStudy (matheducatormcg):

So, assume \[p | x^p+y^p\] and show \[p^2 | x^p+y^p\], directly. Is it trivial?

OpenStudy (zarkon):

it's not too bad i don't think

OpenStudy (zarkon):

start by using fermat on \(x^p\) and \(y^p\)

OpenStudy (matheducatormcg):

oh

OpenStudy (zarkon):

\[x^p\equiv x~ (mod~ p)\] \[y^p\equiv y~ (mod~ p)\]

OpenStudy (zarkon):

\[(x^p+y^p)\equiv x+y~ (mod~ p)\]

OpenStudy (zarkon):

right?

OpenStudy (matheducatormcg):

yes

OpenStudy (matheducatormcg):

raise to the power p?

OpenStudy (zarkon):

\[p|x^p+y^p\] so \[p|x+y\]

OpenStudy (zarkon):

since \(x^p+y^p\) and \(x+y\) are congruent mod \(p\)

OpenStudy (zarkon):

ok?

OpenStudy (matheducatormcg):

yes. thank you. I was over thinking the contradiction approach.

OpenStudy (zarkon):

if \(p|x+y\) then \(x+y=kp\) for some integer \(k\) therefore \(x=kp-y\)

OpenStudy (zarkon):

then look at \[x^p+y^p=(kp-y)^p+y^p\] expand with the binomial theorem

OpenStudy (zarkon):

then use the fact that p is odd to simplify

OpenStudy (matheducatormcg):

thank you